# Transfer Learning for Neutrino Scattering: Domain Adaptation with GANs

### v1

@@ IV Numerical results
To study the effectiveness of transfer learning, we initiated our work by obtaining a baseline (pre-trained) generator and discriminator for inclusive neutrino-carbon scattering events. We trained the model on samples uniformly distributed across different neutrino energies (from $300$ MeV to $10$ Gev), using a training dataset comprising four million events. To test the quality of this model, we generated new data at fixed neutrino energies: $500$ MeV, $1$ GeV, $3$ GeV, $5$ GeV, and $10$ GeV, as well as for uniformly distributed samples across the entire energy spectrum. The values of performance metrics of the baseline GAN model are low and comparable to those reported in our previous paper. Table 1 summarizes the quality of the baseline model.
To evaluate the effectiveness of the transfer learning technique – specifically, to determine how well the baseline model captures universal patterns shared across all processes discussed in this paper – we consider two training scenarios. In particular, we train models using both a large dataset of 100,000 and a smaller one containing 10,000 events.
@@
We begin our discussion with case (A): neutrino scattering on an argon target.
@@
To highlight the differences in optimization between training on 10,000 and 100,000 events, Fig. 5 shows the evolution of performance metrics during training. For the training with 100,000 events, the model’s performance remains stable. It is noteworthy that the best-performing models in the transfer learning framework are reached significantly faster than those trained from scratch. Furthermore, as summarized in Table 2, the transfer learning models consistently achieve lower metric values, indicating better performance. This suggests that when trained from scratch, the network struggles to learn the features of $\nu_{\mu}$–argon scattering from limited data, whereas the transfer learning approach, which builds on prior knowledge of $\nu_{\mu}$–carbon interactions, offers a substantial advantage. Eventually, when the number of samples in the training dataset increases, the quality of the obtained GAN models improves. However, even with a large number of samples, the transfer learning model still outperforms the model trained from scratch.
@@
Figs. 6 and 7 show the event distributions in reconstructed hadronic invariant mass $W$ generated by our GAN models and NuWro for training datasets containing 10,000 and 100,000 events, respectively. This quantity is defined by $W=\sqrt{M_{N}^{2}+2\omega M-Q^{2}}$, with $\omega=E_{\nu}-E_{\mu}$ the energy transferred by the neutrino to the struck nucleus, $M$ the average nucleon mass, and $Q^{2}$ the four-momentum transfer squared $Q^{2}=2E_{\nu}(E_{\mu}-p_{\mu,z})-m_{\mu}^{2}$. The $W$ distributions clearly highlight the quasi-elastic and $\Delta(1232)$ resonance peaks. We show simulations for several neutrino energies: $500$ MeV, $1$, 33, and $5$ GeV. As illustrated, the GAN model trained from scratch struggles to capture the correct structure of the distributions across all energy values. In particular, Fig. 6 highlights the model’s failure to reproduce the peak in the $W$ distribution, whereas the model pre-trained on carbon data achieves a remarkably accurate fit. While using a larger, more informative training dataset improves the performance of the model trained from scratch, it still does not reach the accuracy achieved through transfer learning.
@@
Similarly to case (A), we perform an analogous analysis for case (B), involving $\bar{\nu}_{\mu}$–carbon interactions. The results are summarized in Table 3. Models trained using transfer learning consistently outperform those trained from scratch. When trained on a larger and more informative dataset, model performance improves significantly.
@@
Figs. 8 and 9 show histograms generated by our models and by NuWro for energies of $500$ MeV, and $1$, 33, and $5$ GeV. In contrast to the neutrino–argon case, even the models trained from scratch can capture the resonance structures present in the distributions. Nevertheless, models trained with transfer learning demonstrate superior efficiency and performance.
@@
In the final case (C), we optimize two models, one trained from scratch and the other using transfer learning, using only a small training dataset. Table 4 and Fig. 10 summarize our results. The model trained from scratch struggles to accurately capture the QE peak and performs poorly in the resonance region. In contrast, the model trained via transfer learning performs significantly better, despite being initially trained on data generated by a completely different NuWro model.
@@ V Summary
We have trained GAN models using transfer learning from a pre-trained GAN originally developed on the inclusive CC $\nu_{\mu}$–carbon MC dataset, adapting them to various other scattering scenarios, including inclusive CC $\bar{\nu}_{\mu}$–carbon, $\nu_{\mu}$–argon, and $\nu_{\mu}$–carbon with alternative interaction models. These models were compared against their counterparts trained from scratch. We observed that initializing from the pre-trained model consistently outperforms training from scratch, particularly when datasets have limited statistical significance or provide a poor sampling of the kinematic distributions of the underlying physical processes. This suggests that the pre-trained model captures essential features of neutrino scattering dynamics.
@@
To evaluate the effectiveness of transfer learning more thoroughly, we also trained the same network architecture on datasets with a significantly larger number of events. This demonstrated that the inferior performance of models trained from scratch on smaller datasets is not due to architectural limitations. In this larger-statistics scenario, the model trained from scratch begins to catch up with the transfer-learned model. Nevertheless, the latter continues to show advantages, particularly in regions where the dataset remains a poor representation of the relevant kinematic phase space.
@@
These findings highlight the promising potential of transfer learning in developing more accurate models for (anti)neutrino–nucleus scattering. In particular, one can start from a model trained on MC simulation data and adapt it to more realistic scenarios using comparatively limited data. This approach offers a practical path forward, especially in cases where experimental statistics are sparse. Further investigation in this direction is planned for future work.
Appendix A Loss function
Figure 11: Kinematic distributions for inclusive CC $\nu_{\mu}$-carbon scattering for $E^{\prime}_{\mu}$ and $\theta^{\prime}$ variables computed for 1M events at different neutrino energies, according to NuWro. From left to right and top to bottom distributions for energies $E_{\nu}=$ $0.35$, $0.5$, 11, 33, 55, and $9$ GeV.
Figure 12: Probability distributions from the discriminator for samples generated by NuWro and by the GAN trained with the heuristic loss $L_{H}$ [Eq. (4)] and with our loss $L_{HCE}$ [Eq. (6)], for $E_{\nu}=350$ MeV. Here, NN and $G$ denote samples from NuWro and the GAN, respectively.
Typically, in most applications, the heuristic non-saturating loss
$L_{H}(G)=-\frac{1}{B}\sum_{k=1}^{B}\log[D(G(\mathbf{x}k,E{\nu,k}),E_{\nu,k})]$ (4)
The rationale for using the loss (4) stems from the fact that the maximization of the conventional cross-entropy loss,
$L_{CE}(G)=\frac{1}{B}\sum_{k=1}^{B}\log[1-D(G(\mathbf{x}k,E{\nu,k}),E_{\nu,k})],$ (5)
However, we observe that employing the non-saturating loss in Eq.(4) can lead to uneven performance of the discriminator across different $E_{\nu}$ regions. Specifically, the GAN appears to favor optimization at higher neutrino energies, where the model benefits from smoother kinematic variations across adjacent energy bins. In contrast, at lower energies, slight variations in $E_{\nu}$ produce significant changes in the corresponding kinematic distributions, making training more challenging. The kinematic domains covered by the data for various neutrino energies are illustrated in Fig. 11. At low $E_{\nu}$ energies, $350$ and $500$ MeV, mostly the quasi-elastic contribution is visible. In contrast, at higher energies, resonance contributions also play a significant role. Since higher-energy events dominate our training data, the models being trained first capture the structures associated with these higher energies.
Figure 13: Same as Fig. 12, but for $E_{\nu}=5$ GeV.
Figure 14: Kinematic distributions, pull per bin, and pull distributions for samples from NuWro and GANs trained with the heuristic loss, for $E_{\nu}=500$ MeV (first two rows) and $E_{\nu}=3$ GeV (last two rows).
Figure 15: Same as in Fig. 14 but for model optimized with our loss (5).
For an optimal model of GAN, one can expect the response of the discriminator, which is a probability given by the sigmoid function $Sigmoid(D)={1}/{1+\exp(D)}$, to be distributed around the value $0.5$. Such distributions are shown in Figs. 12 and 13, for low and higher neutrino energy events with $E_{\nu}=350$ MeV and $E_{\nu}=5$ GeV, respectively. These figures display the discriminator distributions for NuWro (“true”) and generated by the GAN distribution of events. The GAN was trained with heuristic loss $L_{H}$ defined in Eq. 4. We observe that the models trained with the heuristic loss exhibit greater dispersion, in low neutrino energy, in their predictions, whereas at higher energies, their distributions are more tightly constrained. Moreover, the lower energy distribution is peaked at a value lower than $0.5$ in contrast to the higher energy distribution. These deficiencies of the discriminator, in low neutrino energies, lead to inefficiency in the optimization of the GAN model.
$L_{HCE}(G)$ $=$ $\frac{1}{B}\sum_{k=1}^{B}\left\{\log\left[1-D(G(\mathbf{x}_{k},E_{\nu,k}),E_{\nu,k})\right]\right$. (6)
$\left.-\log\left[D(G(\mathbf{x}_{k},E_{\nu,k}),E_{\nu,k})\right]\right\},$
which combines the original non-saturating heuristic loss (4) with the maximization term from the conventional cross-entropy (5).
The distributions of discriminator probabilities, obtained with (6), are also shown in Figs. 12 and 13. The pathological shape of the discriminator distribution at low energies disappears, and it is peaked at the value $0.5$.
Additionally, in Fig. 14 and 15 we show 2-dimensional distributions of proxy variables for low and higher neutrino energies ($E_{\nu}=0.5$ and $E_{\nu}=3$ GeV) generated by NuWro and GAN models obtained with heuristic and proposed loss by us, respectively. We observe that at lower neutrino energies, the GAN model trained with heuristic loss fails to reproduce a fraction of events on the boundary of the allowed region.
Figure 16: Dependence on $W$ of the discriminator-predicted probabilities for events generated by GANs trained with the loss functions in Eqs. (4) (left) and (5) (right), for $E_{\nu}=350$ MeV.
Figure 17: Same as in Fig. 16 but for $E_{\nu}=5$ GeV.
Eventually, in Figs. 16 and 17, we show a comparison between $W$-dependence of probability distributions predicted by discriminators for the events generated by the GANs obtained using (4) and (5) loss functions. We observe that the model optimized with the loss proposed in this paper exhibits a distribution concentrated around the probability $0.5$ for both presented neutrino energies.
In contrast, for the model with heuristic loss for low neutrino energy, Fig. 16, besides the events concentrated around probability $0.5$, there are two groups distributed around probability values $0.45$ and $0.55$ for $W$ about $0.85$ and $1.05$ GeV, respectively. These two regions correspond to quasi-elastic and $\Delta$ resonance production contributions.
For higher neutrino energies, Fig. 17, the $W$-distributions for both models are very similar.

### v2

@@ IV Numerical results
To study the TL approach, we began by obtaining a baseline (pre-trained) generator and discriminator for inclusive neutrino-carbon scattering events. We trained the model on samples uniformly distributed across neutrino energies from $300$ MeV to $10$ Gev, using a training dataset with four million events. To test the quality of this model, we generated new data at fixed neutrino energies: $500$ MeV, $1$ GeV, $3$ GeV, $5$ GeV, and $10$ GeV, as well as for uniformly distributed samples across the entire energy spectrum. The values of performance metrics of the baseline GAN model are low and comparable to those reported in our previous paper [23]. Table 1 summarizes the quality of the baseline model.
To evaluate the effectiveness of the TL technique – specifically, to determine how well the baseline model captures universal patterns shared across all processes discussed in this paper – we consider two training scenarios. In particular, we train models using both a large dataset of 100,000 events and a smaller dataset of 10,000 events.
@@
Using a smaller dataset to fine-tune models mimics experimental reality, where data sets are often limited and sparse. Testing the TL approach in this context allows us to evaluate its potential for tuning generative models with real measurements. It’s important to note that experimental data is typically produced in a more complex form, as flux-averaged. In this paper, we present a toy model approach to address this issue, while a more realistic solution is the focus of our current investigations.
@@
We begin our discussion with case (A): neutrino scattering on argon target. To highlight the role of training statistics, we compare models trained on 10,000 and 100,000 events. Fig. 5 shows the evolution of performance metrics during training. For the training with 100,000 events, the model’s performance remains stable. It is noteworthy that the best-performing models in the transfer learning framework are reached significantly faster than those trained from scratch. Furthermore, as summarized in Table 2, the TL models consistently achieve lower metric values, indicating better performance. This suggests that when trained from scratch, the network struggles to learn the features of $\nu_{\mu}$–argon scattering from limited data, whereas the transfer learning approach, which builds on prior knowledge of $\nu_{\mu}$–carbon interactions, offers a substantial advantage. Eventually, when the number of samples in the training dataset increases, the quality of the obtained GAN models improves. However, even with a large number of samples, the transfer learning model still outperforms the model trained from scratch.
Fig. 6 and Figs. S2 and S3 in the supplement show the event distributions in reconstructed hadronic invariant mass $W$ as generated by our GAN models and NuWro for training datasets containing 10,000 and 100,000 events. This quantity is defined by
$W=\sqrt{M_{N}^{2}+2\omega M-Q^{2}},\quad\omega=E_{\nu}-E_{\mu},$ (5)
$M_{N}$ is the average nucleon mass, and the four-momentum transfer squared
$Q^{2}=2E_{\nu}(E_{\mu}-p_{\mu,z})-m_{\mu}^{2},$ (6)
@@
where $p_{\mu,z}$ is the muon momentum component along the neutrino beam direction. The $W$ distributions clearly highlight the QE and $\Delta(1232)$ resonance peaks. In Fig. 6, we show the results from simulations for neutrino energy $500$ MeV, while for neutrino energies: $1$, 33, and $5$ GeV, the results are presented in the supplement. As illustrated, the GAN model trained from scratch struggles to capture the correct structure of the distributions across all energy values. In particular, Fig. 6 highlights the model’s failure to reproduce the peak in the $W$ distribution, whereas the model pre-trained on carbon data achieves a remarkably accurate fit. While using a larger, more informative training dataset improves the performance of the model trained from scratch, it still does not reach the accuracy achieved through TL.
@@
Similarly to the case (A), we perform an analogous analysis for the case (B), involving $\bar{\nu}_{\mu}$–carbon interactions. The results are summarized in Table 3. Models trained using transfer learning consistently outperform those trained from scratch, see Table 3. When trained on a larger and more informative dataset, model performance improves significantly in both scenarios.
@@
Fig. 7 and corresponding figures (Figs. S4 and S5) in the supplement show histograms generated by our models and by NuWro for energies of $500$ MeV, and $1$, 33, 55 GeV, respectively. In contrast to the neutrino–argon case, even the models trained from scratch can capture the resonance structures present in the distributions. Nevertheless, models trained with TL demonstrate superior efficiency and performance.
@@
In the final case (C), we optimize two models, one trained from scratch and the other using TL, using only a small training dataset. Table 4 and Fig. 8 (and Fig. S6) summarize our results. The model trained from scratch struggles to accurately capture the QE peak and performs poorly in the resonance region. In contrast, the model trained via transfer learning performs significantly better, despite being initially trained on data generated by a completely different NuWro model.
As discussed in Sec. III, the event distributions for the four considered processes exhibit significant differences arising from the use of spectral functions versus the LFG model, as well as from the different ways in which neutrinos and antineutrinos interact with nucleons. This is illustrated in Fig. 9, where we show the event distributions as functions of $Q^{2}$ (top panel) and $W$ (bottom panel).
In particular, the antineutrino QE peak is more pronounced, though narrower, than in the case of neutrino–nucleus interactions. The neutrino–carbon and neutrino–argon distributions look similar; however, the QE peak is slightly higher for the scattering off the argon nucleus. In contrast, the alternative nuclear model based on the LFG approximation, combined with a different resonance description, produces event distributions with visibly different shapes and notably different strengths of both the QE and the $\Delta(1232)$ resonance peaks.
Figure 9: Event distributions in $Q^{2}$ and $W$ for $E_{\nu}=1$ GeV were obtained from our GAN models for neutrino-carbon, neutrino-argon, antineutrino-carbon, and an alternative model of neutrino-carbon scattering processes. The models were trained using 10,000 events and transfer learning.
@@ V Summary
We have trained GAN models using transfer learning from a pre-trained GAN originally developed on the inclusive CC $\nu_{\mu}$-carbon scattering MC dataset, adapting them to various other scattering scenarios, including inclusive CC $\bar{\nu}_{\mu}$–carbon, $\nu_{\mu}$-argon, and $\nu_{\mu}$–carbon with alternative interaction models. These models were compared against their counterparts trained from scratch. We observed that initializing from the pre-trained model consistently outperforms training from scratch, particularly when datasets are small or poorly sample the kinematic distributions of the underlying physical processes. This suggests that the pre-trained model captures essential features of neutrino scattering dynamics.
@@
To more thoroughly evaluate the effectiveness of transfer learning, we also trained the same network architecture on datasets with a significantly larger number of events. This demonstrated that the inferior performance of models trained from scratch on smaller datasets is not due to architectural limitations. In this larger-statistics scenario, the model trained from scratch begins to catch up with the transfer-learned model. Nevertheless, the latter continues to show some advantages, particularly in regions where the dataset remains a poor representation of the relevant kinematic phase space.
From a technical standpoint, the frozen initial layers of the generator primarily reshape the latent Gaussian distribution into a generic lepton-kinematic manifold that should be rather universal for various neutrino-induced processes. The subsequent unfrozen layers adapt this manifold to process-specific features. Therefore, the first block is effectively universal for all neutrino and antineutrino scattering tasks, while fine-tuning later layers successfully transfers the model to the target domain. In the case of the discriminator, which takes lepton kinematics as input, the first frozen block is also rather universal and extracts/processes basic features that subsequent blocks transform. These blocks are adapted to a particular type of process. The fact that TL from neutrino-carbon to neutrino–argon and antineutrino-carbon succeeds demonstrates that the GAN primarily encodes universal kinematic constraints and the overall QE and $\Delta$-resonance pattern, while target-dependent spectral-function details are corrected by fine-tuning in subsequent layers.
There are several physical arguments supporting our approach. First: in all the discussed processes, the underlying nuclear physics is the same. The carbon and argon targets differ in nuclear structure, but, as we showed in our previous paper, the TL approach can be less effective when the differences are radical, as in the case of 3He and 12C targets. However, even in this case, the TL scenario in full-tuning mode still works efficiently. Indeed, when one looks at the QE and $\Delta$ peaks for carbon and argon, they differ but not dramatically. Hence, the carbon-scattering model can be considered a good a posteriori model for argon. A similar argument works for neutrino and antineutrino models. In this case, the target is the same, but the neutrino and antineutrino probe nuclear structure, which is reflected in some differences between the resulting distributions. However, minor tuning allows us to obtain one model from the other.
@@
Our findings highlight the promising potential of transfer learning in developing accurate models for (anti)neutrino–nucleus scattering. In particular, one can start from a model trained on MC simulation data and adapt it to more realistic scenarios using comparatively limited data. This approach offers a practical path forward, especially when experimental statistics are sparse. Further investigation in this direction is planned for future work, as well as extending the method to full final-state hadrons, which is part of the ongoing work.
@@ Acknowledgements.
Supplemental Material
S1 Event distributions
Figure S1: Test data samples $(E_{\mu},\theta)$ distributions for $E_{\nu}=500$ MeV. On the left, we have the original $\nu_{\mu}$-carbon scattering; on the right, the $\nu_{\mu}$-argon scattering.
In the main paper, we show the distribution of events, in proxy variables, for all the processes we study. Here, in Fig. S1, we additionally present the analogical distributions but in physical variables $(E_{\mu},\theta)$ for neutrino-carbon and neutrino-argon scattering. Spectral functions describe the carbon and argon nuclei from Refs. [33] and [38, 39], respectively.
S2 Reconstructed hadronic invariant mass $W$-distributions
This section contains the distribution of events generated by GANs for neutrino-argon scattering for models trained on 10,000 (Fig. S2) and 100,000 (Fig. S3) events, respectively. Similarly in Figs. S4 and S5 show the distribution of events generated by the GANs trained on 10,000 and 100,000 events for antineutrino-carbon scattering. Eventually, the Fig. S6 shows histograms for events generated by the model of neutrino-carbon trajectory on 10,000 events obtained from NuWro using an alternative interaction model.
Figure S2: Reconstructed hadronic invariant mass $W$-distributions, for inclusive CC $\nu_{\mu}$-argon scattering, generated by NuWro and GAN model trained from scratch and using TL, for the optimization with 10k events. We present the results for $E_{\nu}=500$ MeV (top left), 11 GeV (top right), 33 GeV (bottom left), and 55 GeV (bottom right).
Figure S3: Same as in Fig. S2 but for $C$ $\nu_{\mu}$-argon scattering and the training dataset with 100,000 events.
Figure S4: Same as in Fig. S2 but for $C$ $\bar{\nu}_{\mu}$-carbon scattering and the training dataset with 10,000 events.
Figure S5: Same as in Fig. S2 but for $C$ $\bar{\nu}_{\mu}$-carbon scattering and the training dataset with 100,000 events.
Figure S6: Same as in Fig. S2 but for $C$ $\nu_{\mu}$-carbon (alternative interaction) scattering and the training dataset with 10,000 events.
S3 Loss function
Figure S7: Kinematic distributions for inclusive CC $\nu_{\mu}$-carbon scattering for $E^{\prime}_{\mu}$ and $\theta^{\prime}$ variables computed for 1M events at different neutrino energies, according to NuWro. From left to right and top to bottom distributions for energies $E_{\nu}=$ $0.35$, $0.5$, 11, 33, 55, and 99 GeV.
Figure S8: Probability distributions from the discriminator for samples generated by NuWro and by the GAN trained with the heuristic loss $L_{H}$ [Eq. S1] and with our loss $L_{HCE}$ [Eq. S3], for $E_{\nu}=350$ MeV. Here, NN and $G$ denote samples from NuWro and the GAN, respectively.
Typically, in most applications, the heuristic non-saturating loss [15]
$L_{H}(G)=-\frac{1}{B}\sum_{k=1}^{B}\log[D(G(\mathbf{x}k,E{\nu,k}),E_{\nu,k})]$ (S1)
The rationale for using the loss (S1) stems from the fact that the maximization of the conventional cross-entropy loss,
$L_{CE}(G)=\frac{1}{B}\sum_{k=1}^{B}\log[1-D(G(\mathbf{x}k,E{\nu,k}),E_{\nu,k})],$ (S2)
However, we observe that employing the non-saturating loss in Eq. S1 can lead to uneven performance of the discriminator across different $E_{\nu}$ regions. Specifically, the GAN appears to favor optimization at higher neutrino energies, where the model benefits from smoother kinematic variations across adjacent energy bins. In contrast, at lower energies, slight variations in $E_{\nu}$ produce significant changes in the corresponding kinematic distributions, making training more challenging. The kinematic domains covered by the data for various neutrino energies are illustrated in Fig. S7. At low $E_{\nu}$ energies, $350$ and $500$ MeV, mostly the quasi-elastic contribution is visible. In contrast, at higher energies, resonance contributions also play a significant role. Since higher-energy events dominate our training data, the models being trained first capture the structures associated with these higher energies.
Figure S9: Same as Fig. S8, but for $E_{\nu}=5$ GeV.
Figure S10: Kinematic distributions, pull per bin, and pull distributions for samples from NuWro and GANs trained with the heuristic loss, for $E_{\nu}=500$ MeV (first two rows) and $E_{\nu}=3$ GeV (last two rows).
Figure S11: Same as in Fig. S10 but for model optimized with our loss (S2).
For an optimal model of GAN, one can expect the response of the discriminator, which is a probability given by the sigmoid function $Sigmoid(D)={1}/{1+\exp(D)}$, to be distributed around the value $0.5$. Such distributions are shown in Figs. S8 and S9, for low and higher neutrino energy events with $E_{\nu}=350$ MeV and $E_{\nu}=5$ GeV, respectively. These figures display the discriminator distributions for NuWro (“true”) and generated by the GAN distribution of events. The GAN was trained with heuristic loss $L_{H}$ defined in Eq. S1. We observe that the models trained with the heuristic loss exhibit greater dispersion, in low neutrino energy, in their predictions, whereas at higher energies, their distributions are more tightly constrained. Moreover, the lower energy distribution is peaked at a value lower than $0.5$ in contrast to the higher energy distribution. These deficiencies of the discriminator, in low neutrino energies, lead to inefficiency in the optimization of the GAN model.
$L_{HCE}(G)$ $=$ $\frac{1}{B}\sum_{k=1}^{B}\left\{\log\left[1-D(G(\mathbf{x}_{k},E_{\nu,k}),E_{\nu,k})\right]-\log\left[D(G(\mathbf{x}_{k},E_{\nu,k}),E_{\nu,k})\right]\right\},$ (S3)
which combines the original non-saturating heuristic loss (S1) with the maximization term from the conventional cross-entropy (S2).
The distributions of discriminator probabilities, obtained with (S3), are also shown in Figs. S8 and S9. The pathological shape of the discriminator distribution at low energies disappears, and it peaks at $0.5$.
Figure S12: Dependence on $W$ of the discriminator-predicted probabilities for events generated by GANs trained with the loss functions in Eqs. (S1) (left) and (S2) (right), for $E_{\nu}=350$ MeV.
Figure S13: Same as in Fig. S12 but for $E_{\nu}=5$ GeV.
Additionally, in Figs. S10 and S11 we show 2-dimensional distributions of proxy variables for low and higher neutrino energies ($E_{\nu}=0.5$ and $E_{\nu}=3$ GeV) generated by NuWro and GAN models obtained with heuristic and proposed loss by us, respectively. We observe that at lower neutrino energies, the GAN model trained with heuristic loss fails to reproduce a fraction of events on the boundary of the allowed region.
Eventually, in Figs. S12 and S13, we show a comparison between $W$-dependence of probability distributions predicted by discriminators for the events generated by the GANs obtained using (S1) and (S2) loss functions. We observe that the model optimized with the loss proposed in this paper exhibits a distribution concentrated around the probability $0.5$ for both presented neutrino energies.
In contrast, for the model with heuristic loss for low neutrino energy, Fig. S12, besides the events concentrated around probability $0.5$, there are two groups distributed around probability values $0.45$ and $0.55$ for $W$ about $0.85$ and $1.05$ GeV, respectively. These two regions correspond to quasi-elastic and $\Delta$ resonance production contributions.
For higher neutrino energies, Fig. S13, the $W$-distributions for both models are very similar.